8. A 30 kg object, initially at rest, slides 10 m down a 20° ramp at 1. 8 m/s2. a. How long does it take the object to get to the bottom of the ramp? b. what is the coefficient of friction between the object and the ramp?
5Kg) (1.5m/s)m m, :0.45 d) The students are now asked to consider the kinetic energy changes in an inelastic collision, specifically whether the initial values of one of the physical quantities affect the fraction of mechanical energy dissipated in the collision. How could you modify the experiment to investigate this question? Be sure to explicitly describe the calculations you would make, specifying all equations you would use (but do not actually do any algebra or arithmetic).
Answer #1

here,

8)

mass of object , m = 30 kg

s = 10 m

theta = 20 degree

accelration , a = 1.8 m/s^2

a)

let the time taken be t

s = v0 * t + 0.5 * a * t^2

10 = 0 + 0.5 * 1.8 * t^2

solving for t

t = 3.33 s

the time taken is 3.33 s


b)

let the coefficient of friction be uk

accelration , a = net force /effective mass

a = ( m * g * sin(theta) - uk * m * g * cos(theta))/m

1.8 = 9.81 * (sin(20) - uk * cos(20))

solving for uk

uk = 0.17